Nimbus, Inc., makes brooms and then sells them to customers. Here is the relationship between the number of workers and Nimbus's output in a given day:


Workers:

  1. 0 worker
  2. 1 worker
  3. 2 workers
  4. 3 workers
  5. 4 workers
  6. 5 workers
  7. 6 workers
  8. 7 workers


Output:

  1. 0 broom
  2. 20 brooms
  3. 50 brooms
  4. 90 brooms
  5. 120 brooms
  6. 140 brooms
  7. 150 brooms
  8. 155 brooms


a.)Marginal product / b.)Total cost / c.)Average Total cost / d.)Marginal cost:


  1. ?
  2. ?
  3. ?
  4. ?


a.)  Construct the marginal-cost and average-total-cost curves for Nimbus. Explain diminishing marginal product and explain when does Nimbus experience diminishing marginal product using the cost curves below.


1
Expert's answer
2021-08-17T16:57:00-0400

a)

"Marginal\space product= \frac{Change\space in\space output}{Change\space in\space worker}"




Marginal product initially increases till 3 units of worker and after this marginal product starts to decrease.

The reason for this pattern is due to return to a factor and law of variable proportion.


b.

Worker cost= 100

Fixed cost= 200

Total cost= 200+100(number of worker)





c.

"ATC = \frac{ TC}{output}"





ATC initially decreases and after 5 unit of worker it starts to rise. Here ATC is U-shaped



d.

"Marginal\space cost= \frac{Change\space in\space TC}{Change\space in\space Q}"




MC initially decreases and after 90 units of output it starts to rise. Here MC is U-shaped


a)






Diminishing marginal returns holds that increase in variable input units holding a certain amount of fixed inputs, the output initially grow at a faster rate, then at a steady rate, but ultimately, it will grow at a declining rate.


 Nimbus experience diminishing marginal product  starting from adding the fourth worker.